OpenStudy (anonymous):

−5(x − 2) = 15 solve please

OpenStudy (anonymous):

First add 5 to both sides of the equation

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

Divide both sides by -5 first*

OpenStudy (anonymous):

ok what else(: & thank you

OpenStudy (anonymous):

once you do that you have x-2=-3 now add 2 to both sides

OpenStudy (anonymous):

so its -1 ?

OpenStudy (anonymous):

Yes x=-1

OpenStudy (anonymous):

THANK YOU SO MUCH ! :D can u help me with another problem & i promise i wont bother anymore

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

2x + 36 + 40 = 180

OpenStudy (anonymous):

Simplify the left side first, add the numbers

OpenStudy (anonymous):

how im so confused :( im so stupid

OpenStudy (anonymous):

whats 36+40?

OpenStudy (anonymous):

76

OpenStudy (anonymous):

right, so now you have 2x+76=180 now subtract 76 from both sides

OpenStudy (anonymous):

104?

OpenStudy (anonymous):

is 52 the answer?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

yay thank you